OpenStudy (anonymous):

Use this polynomial, f(x)=3x^4-11x^3+10x-4 To find all possible rational zeros, to describe the roots of the polynomial, find an upper bound, find a lower bound, state all the solutions over the set of complex numbers, factor the polynomial completely over the set pf complex numbers, and prove the linear factors produce the polynomial. If I have someone just to guided me would be Helpful. WILL FAN AND MEDAL.

OpenStudy (anonymous):

I know its a lot but I do need help.

OpenStudy (anonymous):

i can help with the first part

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

all possible rational zeros, numbers that divide 4 (the constant)

OpenStudy (anonymous):

oops i didn't read it carefully scratch that

OpenStudy (anonymous):

all fractions where the numerator divides 4 (the constant) and the denominator divides 3 (the leading coefficient)

OpenStudy (anonymous):

the list is kind of long \[\pm1,\pm2,\pm4\] and then all of those over \(3\)

OpenStudy (anonymous):

You use those numbers because those are factrs of +/- 4

OpenStudy (anonymous):

over 3 cause 3 is leading coefficiannt

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so also include\[\pm\frac{1}{3},\pm\frac{2}{3},\pm\frac{4}{3}\] that is all

OpenStudy (anonymous):

Oh duh.. I knew that.. So, those would be rational zeros,?

OpenStudy (anonymous):

OKAY. Thanks, can you explain Decartes Rule by any chance?

OpenStudy (anonymous):

oh no it is a cubic polynomial, it has at most 3 real zeros that is a list of POSSIBLE RATIONAL zeros

OpenStudy (anonymous):

ok i mis read it again it is a polynomial of degree 4 it has at most 4 real zeros

OpenStudy (anonymous):

yes we can look at descartes rule of sine

OpenStudy (anonymous):

I knew it had at most 4 cause its quadtratic in its exponent :) And sine or sign?

OpenStudy (anonymous):

it is not a "quadratic" i know quad is often used to mean "4" like in "quadrophonic" but a quadratic has degree 2 this has degree 4 don't say it is quadratic

OpenStudy (anonymous):

descartes rule of sign

OpenStudy (anonymous):

Oh. Yeah, Its a weird name but yes. THANKS for correcting me.

OpenStudy (anonymous):

\[ f(x)=3x^4-11x^3+10x-4\] has as 3 changes in sign of the coefficients

OpenStudy (anonymous):

Because of the '-' and '+' changing 3 times?

OpenStudy (anonymous):

from \(+3\) to \(-11\) is one change from \(-11\) to \(+10\) is another from \(+10\) to \(-4\) is the third

OpenStudy (anonymous):

Gotcha :)

OpenStudy (anonymous):

that tells you it has 3 positive zeros or 1 (you count down by twos, and since it cannot have \(-1\) positive zeros it either has 3 or 1

OpenStudy (anonymous):

i.e. it MUST have one positive zero, maybe 3

OpenStudy (anonymous):

now for the negative zeros, look at the number of changes in sign of \(f(-x\)

OpenStudy (anonymous):

\[f(x)=3x^4-11x^3+10x-4\] \[f(-x)=3(-x)^4-11(-x)^3+10(-x)-4\]\[=3x^4+11x^3-10x-4\]

OpenStudy (anonymous):

how many changes in sign do you see ?

OpenStudy (anonymous):

+3 + +11 is first +11 - 10 is two -10 - 4 is third, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

or just 2?

OpenStudy (anonymous):

CHANGES in sign from \(+\) to \(-\) or \(-\) to \(+\)

OpenStudy (anonymous):

yeah it only does that twice at the beginning.

OpenStudy (anonymous):

i mean once.. right?

OpenStudy (anonymous):

Or am i over thinking now

OpenStudy (anonymous):

look carefully yes once

OpenStudy (anonymous):

okay now I see

OpenStudy (anonymous):

from \(+11\) to \(-3\)

OpenStudy (anonymous):

ts wanting to see changes in addition to subtraction or vice versa

OpenStudy (anonymous):

ok not really from \(+11\) to \(-10\)

OpenStudy (anonymous):

and in how many times, to tell us the roots we may have

OpenStudy (anonymous):

for \(f(-x)\) it counts the NEGATIVE roots

OpenStudy (anonymous):

So -11+10 only counts?

OpenStudy (anonymous):

Ahh. and -11 is the first negative

OpenStudy (anonymous):

So we start from there

OpenStudy (anonymous):

are you looking at \(f(x)\) or at \(f(-x)\)?

OpenStudy (anonymous):

Oh I was looking at the wrong one, I am sorry. Scrolled up too far. I am so sorry

OpenStudy (anonymous):

So its looking for the first subtraction part?

OpenStudy (anonymous):

\[f(x)=3x^4-11x^3+10x-4\] three changes \[f(-x)=3x^4+11x^3-10x-4\] one change

OpenStudy (anonymous):

For negative roots?

OpenStudy (anonymous):

one change for \(f(-x)\) means one negative root

OpenStudy (anonymous):

three changes for \(f(x)\) means three or one positive roots

OpenStudy (anonymous):

Okay, I see. I had to process that. Now, is it always true for f(-x) to have one negative root, or can there also be 3 negative rots, depending on the problem?

OpenStudy (anonymous):

there could be as many as there are changes in sign

OpenStudy (anonymous):

That one just happened to have 1.

OpenStudy (anonymous):

for negative, anyways

OpenStudy (anonymous):

and you count down by twos so for example if there were 5 changes in sign of \(f(-x)\) there could be 5,3 or 1 negative root

OpenStudy (anonymous):

same with positive roots when looking at the changes of sign of \(f(x)\)

OpenStudy (anonymous):

I see, Thats why there is 3 or 1 positive for the f(x) positive

OpenStudy (anonymous):

in your example only one change in sign of \(f(-x)\) so one negative root

OpenStudy (anonymous):

I see now. That rule isnt too bad!

OpenStudy (anonymous):

got stuck sorry

OpenStudy (anonymous):

now to find the zeros i guess you have to factor this

OpenStudy (anonymous):

by some miracle it factors as \[(3 x-2) (x+1) (x^2-4 x+2)\] so the zeros are easy to find

OpenStudy (anonymous):

@satellite73 would that be considered the solutions?

OpenStudy (anonymous):

Whaever makes those =0